OpenStudy (anonymous):

When do you feel weightless or lightweight? I mean, for example, if you are in an elevator accelerating upward, what makes you feel the weightgain? Is it normal reaction? Or when elevator accelerates downward, what makes u feel the weightloss? I know that the weighing scale shows change in weight due to corresponding change in normal reaction but what makes YOU feel it? If you jump, do you feel weightless while rising also or only while falling.

OpenStudy (anonymous):

do you understand the q?

OpenStudy (anonymous):

i can give you an answer that u will be very satisfied with but he will give you an expalnation that will remain forever in ur head! http://ocw.mit.edu/courses/physics/8-01-physics-i-classical-mechanics-fall-1999/video-lectures/lecture-7/ start at 26:15

OpenStudy (experimentx):

weight is a force, usually force (if unbalanced) is compensated by acceleration .... equivalently it can be said that effect of acceleration will result force.

OpenStudy (anonymous):

i would have loved to see MIT videos but the internet connection is INSANELY slow

OpenStudy (anonymous):

@experimentX ??

OpenStudy (experimentx):

that's the nature of force.

OpenStudy (anonymous):

lets start with this If you jump, do you feel weightless while rising also or only while falling?

OpenStudy (experimentx):

if you jump you will feel more weight(at that instant of jump), if you fall, you will feel weightlessness.

OpenStudy (anonymous):

when u r only under gravity's influence(free fall===falling under acc of 9.8 oONLY u are weightless the raeson being that......

OpenStudy (anonymous):

u have to consider that u have a weighing scale beneath u as u fall it will raed the reaction force given by you it will be directed up mostly as u fall down so this raection i wish to acll as R when u are going bin alift up, so R shud be more mg is act6ing down non-negotiable the NET force being up so R-mg=ma so wat is R when u are falling down (free fall) oir wen u are going down a lift?

OpenStudy (anonymous):

yep 0 in freefall

OpenStudy (anonymous):

down a lift?

OpenStudy (anonymous):

less than that at constant speed or rest but greater than 0, right?

OpenStudy (anonymous):

assuming lift doesnt go down at g

OpenStudy (anonymous):

yes! why can u rite down the equation pertaining to it? the reaction R?

OpenStudy (anonymous):

yeah!

OpenStudy (ujjwal):

knowledge of micro gravity can prove to be helpful if you want a deeper explanation.. Google it

OpenStudy (anonymous):

k im back net force upward R R-mg=ma R=ma+mg=m(-g)+mg=0

OpenStudy (anonymous):

no u got it wrong....the R normal force upward is more as we do not take g as -9.8 or-10 the raeson being th that when u reverse the sign writing -mg it is beautifully understood that u ahve accounted for the sign

OpenStudy (anonymous):

what did i do wrong this is for free fall no?

OpenStudy (anonymous):

First note that weight means the force due to gravity. When you stand on a scale, gravity is pulling down on you and the scale and your weight is measured. But when you are standing on scale and say falling off a cliff, both the scale and you are being pulled down by gravity equally and therefore because you are not pushing down on the scale and the scale is pushing on you it would read 0 pounds (or kg). So when you are going down in the elevator the reading decreases because the scale is falling partially due to gravity and so your weight is minimized. If it was going up, however, the scale would show a higher weight because both you and the scale are opposing gravity and pushing on each other more.